# Nonasymptotic Analysis of Accelerated Methods With Inexact Oracle Under Absolute Error Bound

### v1

@@ A.1 Proof of composition optimization
$\Pr\left(\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\geq\epsilon\left(\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)\right)$
$\leq$ $\frac{1}{\epsilon}\mathbb{E}\left[\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\right]/\left(\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)$
@@
~$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla f(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla f(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq\epsilon\left(\frac{2C_{f}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{f}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)$
@@
A.2 Convergence rate for inexact generalized FGM
Algorithm 6 Inexact Generalized Fast Gradient Method (iGFGM)
0:  stepsize parameter $\{\lambda_{k}\},\lambda_{k}\in[0,1]$, $\mathbf{y}_{0}=\mathbf{z}_{0}=\mathbf{x}_{0}\in\mathbb{R}^{d}$, $A_{0}=\alpha_{0}=1$.
1:  for $k=0,\cdots,K-1$ do
2:     $\mathbf{y}_{k+1}=\mathbf{x}_{k}-\frac{1}{L}\tilde{\nabla}f(\mathbf{x}_{k})$
3:     $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{1}{L}\alpha_{k}\tilde{\nabla}f(\mathbf{x}_{k})$
4:     $A_{k+1}=A_{k}+\alpha_{k+1}$ with $\alpha_{k+1}=\frac{\lambda_{k+1}+\sqrt{4\lambda_{k+1}A_{k}+\lambda_{k+1}^{2}}}{2}$
5:     $\mathbf{x}_{k+1}=(1-\frac{\alpha_{k+1}}{A_{k+1}})\mathbf{y}_{k+1}+\frac{\alpha_{k+1}}{A_{k+1}}\mathbf{z}_{k+1}$
6:  end for
@@ A.2 Convergence rate for inexact generalized FGM
$A_{K}(f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2})$
$\leq$ $-\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rangle-\sum_{k=0}^{K}(A_{k}-A_{k-1})\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{*}-\mathbf{x}_{k}\right\rangle$
$-\frac{1}{2L}\sum_{k=0}^{K-1}A_{k}\left\|\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{1}{2L}\sum_{k=0}^{K}(A_{k}-A_{k-1})\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{A_{K}}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
$=$ $\sum_{k=0}^{K}\alpha_{k}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}(A_{k+1}\theta_{k+1,i}-A_{k}\theta_{k,i})\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k+1}\theta_{k+1,k}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}A_{k}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
$+\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
@@
$=$ $\sum_{k=0}^{K}\alpha_{k}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}A_{k}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
@@
$-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}(\alpha_{k}\alpha_{k+1}+A_{k})\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
$=$ $-\frac{1}{\tau}\left\|\tau(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\tau\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K}\left(\frac{\alpha_{k}^{2}}{4\tau}-\frac{A_{k}}{L}\right)\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
@@
$+\frac{1}{2\tau}\sum_{k=0}^{K}\sum_{i=0}^{k-1}\alpha_{k}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{i})\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
@@
$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\alpha_{k}\alpha_{k+1}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
$=$ $-\frac{2}{L}\left\|\frac{L}{2}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\sum_{k=0}^{K}\frac{2A_{k}-\alpha_{k}^{2}}{2L}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
@@
$+\frac{1}{L}\sum_{k=1}^{K}\alpha_{k}\alpha_{k-1}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k-1})\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle$
@@
$-\frac{1}{L}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\alpha_{k}\alpha_{k+1}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
$=$ $-\frac{2}{L}\left\|\frac{L}{2}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\sum_{k=0}^{K}\frac{2A_{k}-\alpha_{k}^{2}}{2L}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
@@
$-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}(\alpha_{k}\alpha_{k+1}+A_{k})\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle$
$=$ $-\frac{2}{L}\left\|\frac{L}{2}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\alpha_{k}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\frac{2A_{0}-\alpha_{0}^{2}}{2L}\left\|\nabla f(\mathbf{x}_{0})\right\|^{2}$
$-\sum_{k=0}^{K-1}\frac{1}{2L(2A_{k+1}-\alpha_{k+1}^{2})}\left\|(2A_{k+1}-\alpha_{k+1}^{2})\nabla f(\mathbf{x}_{k+1})+\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\mathbf{e}_{i}+(\alpha_{k}\alpha_{k+1}+A_{k})\mathbf{e}_{k}\right\|^{2}$
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}^{2}\alpha_{i}^{2}}{2L(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{i}\right\|^{2}+\sum_{k=0}^{K-1}\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2L(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
@@
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=0}^{i-1}\frac{\alpha_{k+1}^{2}\alpha_{i}\alpha_{j}}{L(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{i},\mathbf{e}_{j}\right\rangle+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k}\alpha_{k+1}+A_{k})}{L(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{i},\mathbf{e}_{k}\right\rangle$
$\leq$ $\frac{L}{2}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K-1}\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2L(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{2L(2A_{j+1}-\alpha_{j+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k}\alpha_{k+1}+A_{k})}{L(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{L(2A_{j+1}-\alpha_{j+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle.$
@@
$\hat{P}_{k,k}\triangleq$ $\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2L(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{2L(2A_{j+1}-\alpha_{j+1}^{2})}$
$\hat{P}_{k,i}\triangleq$ $\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k+1}\alpha_{k}+A_{k})}{L(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{L(2A_{j+1}-\alpha_{j+1}^{2})},$
@@
$=$ $\sum_{k=0}^{K-1}\left(\frac{A_{k}^{2}(1+\alpha_{k+1})}{2L(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K}\frac{\alpha_{k}A_{i-1}\alpha_{i}(1+\alpha_{i})}{2L(2A_{i}-\alpha_{i}^{2})}\right)\left\|\mathbf{e}_{k}\right\|^{2}.$
Dividing both sides of the inequality by $A_{K}$, the final result is obtained. ∎

### v2

@@ A.1 Details on motivational examples
$\Pr\left(\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}\geq\epsilon\left(\frac{2C_{h}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)\right)$
$\leq$ $\frac{1}{\epsilon}\mathbb{E}\left[\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}\right]/\left(\frac{2C_{h}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)$
@@
$\left\|\tilde{\nabla}\mathbf{g}(\mathbf{x})\nabla h(\tilde{\mathbf{g}}(\mathbf{x}))-\nabla\mathbf{g}(\mathbf{x})\nabla h(\mathbf{g}(\mathbf{x}))\right\|^{2}\leq\epsilon\left(\frac{2C_{h}^{2}}{\left|\mathcal{S}_{\nabla}\right|}\sigma_{\nabla}^{2}+\frac{2C_{g}^{2}L_{h}^{2}}{\left|\mathcal{S}_{\mathbf{g}}\right|}\sigma_{\mathbf{g}}^{2}\right)$
@@
A.2 Convergence bound of iGFGM
@@
$f(\mathbf{x}_{K})-f_{*}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
$\leq$ $-\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rangle-\sum_{k=0}^{K}\frac{A_{k}-A_{k-1}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{*}-\mathbf{x}_{k}\right\rangle$
$-\frac{1}{2L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\|^{2}-\frac{1}{2L}\sum_{k=0}^{K}\frac{A_{k}-A_{k-1}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}-\frac{1}{2L}\left\|\nabla f(\mathbf{x}_{K})\right\|^{2}$
@@
$=$ $\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{0}-\mathbf{x}_{*}\right\rangle-\frac{1}{L}\sum_{k=0}^{K}\frac{A_{k}}{A_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k+1})\right\rangle$
@@
$-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
@@
$-\frac{1}{L}\sum_{k=0}^{K-1}\frac{\alpha_{k}\alpha_{k+1}+A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
$=$ $-\frac{1}{\tau}\left\|\tau(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}+\tau\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K}\left(\frac{\alpha_{k}^{2}}{4\tau A_{K}^{2}}-\frac{A_{k}}{LA_{K}}\right)\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
@@
$+\frac{1}{2\tau}\sum_{k=0}^{K}\sum_{i=0}^{k-1}\frac{\alpha_{k}\alpha_{i}}{A_{K}^{2}}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{i})\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{i})+\mathbf{e}_{i}\right\rangle$
@@
$-\frac{1}{L}\sum_{k=0}^{K-1}\frac{A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\frac{\alpha_{k}\alpha_{k+1}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
$=$ $-\frac{2A_{K}}{L}\left\|\frac{L}{2A_{K}}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2A_{K}}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\sum_{k=0}^{K}\frac{2A_{k}-\alpha_{k}^{2}}{2LA_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
@@
$+\frac{1}{L}\sum_{k=1}^{K}\frac{\alpha_{k}\alpha_{k-1}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k}),\nabla f(\mathbf{x}_{k-1})\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle$
@@
$-\frac{1}{LA_{K}}\sum_{k=0}^{K-1}A_{k}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\frac{\alpha_{k}\alpha_{k+1}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\nabla f(\mathbf{x}_{k})+\mathbf{e}_{k}\right\rangle$
$=$ $-\frac{2A_{K}}{L}\left\|\frac{L}{2A_{K}}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2A_{K}}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\sum_{k=0}^{K}\frac{2A_{k}-\alpha_{k}^{2}}{2LA_{K}}\left\|\nabla f(\mathbf{x}_{k})\right\|^{2}$
@@
$-\frac{1}{L}\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{i}\right\rangle-\frac{1}{L}\sum_{k=0}^{K-1}\frac{\alpha_{k}\alpha_{k+1}+A_{k}}{A_{K}}\left\langle\nabla f(\mathbf{x}_{k+1}),\mathbf{e}_{k}\right\rangle$
$=$ $-\frac{2A_{K}}{L}\left\|\frac{L}{2A_{K}}(\mathbf{x}_{0}-\mathbf{x}_{*})-\frac{1}{2}\sum_{k=0}^{K}\frac{\alpha_{k}}{A_{K}}\nabla f(\mathbf{x}_{k})\right\|^{2}+\frac{L}{2A_{K}}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}-\frac{2A_{0}-\alpha_{0}^{2}}{2LA_{K}}\left\|\nabla f(\mathbf{x}_{0})\right\|^{2}$
$-\sum_{k=0}^{K-1}\frac{1}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\|(2A_{k+1}-\alpha_{k+1}^{2})\nabla f(\mathbf{x}_{k+1})+\sum_{i=0}^{k-1}\alpha_{k+1}\alpha_{i}\mathbf{e}_{i}+(\alpha_{k}\alpha_{k+1}+A_{k})\mathbf{e}_{k}\right\|^{2}$
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}^{2}\alpha_{i}^{2}}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{i}\right\|^{2}+\sum_{k=0}^{K-1}\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=0}^{i-1}\frac{\alpha_{k+1}^{2}\alpha_{i}\alpha_{j}}{LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{i},\mathbf{e}_{j}\right\rangle+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k}\alpha_{k+1}+A_{k})}{LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{i},\mathbf{e}_{k}\right\rangle$
$\leq$ $\frac{L}{2A_{K}}\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|^{2}+\sum_{k=0}^{K-1}\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}+\sum_{k=0}^{K-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{2LA_{K}(2A_{j+1}-\alpha_{j+1}^{2})}\left\|\mathbf{e}_{k}\right\|^{2}$
$+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k}\alpha_{k+1}+A_{k})}{LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle+\sum_{k=0}^{K-1}\sum_{i=0}^{k-1}\sum_{j=k+1}^{K-1}\frac{\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{LA_{K}(2A_{j+1}-\alpha_{j+1}^{2})}\left\langle\mathbf{e}_{k},\mathbf{e}_{i}\right\rangle.$
@@
$\hat{P}_{k,k}\triangleq$ $\frac{(\alpha_{k+1}\alpha_{k}+A_{k})^{2}}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{k}^{2}\alpha_{j+1}^{2}}{2LA_{K}(2A_{j+1}-\alpha_{j+1}^{2})}$
$\hat{P}_{k,i}\triangleq$ $\frac{\alpha_{k+1}\alpha_{i}(\alpha_{k+1}\alpha_{k}+A_{k})}{LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{j=k+1}^{K-1}\frac{\alpha_{j+1}^{2}\alpha_{k}\alpha_{i}}{LA_{K}(2A_{j+1}-\alpha_{j+1}^{2})},$
@@
$=$ $\sum_{k=0}^{K-1}\left(\frac{A_{k}^{2}(1+\alpha_{k+1})}{2LA_{K}(2A_{k+1}-\alpha_{k+1}^{2})}+\sum_{i=k+1}^{K}\frac{\alpha_{k}A_{i-1}\alpha_{i}(1+\alpha_{i})}{2LA_{K}(2A_{i}-\alpha_{i}^{2})}\right)\left\|\mathbf{e}_{k}\right\|^{2}.$